OpenStudy (anonymous):

What polynomial has roots of -4, 2, and 5 ? a.) x3 - x2 - 22x + 40 b.) x3 + x2 - 22x - 40 c.) x3 + 3x2 - 18x - 40 d.) x3 - 3x2 - 18x + 40

OpenStudy (campbell_st):

use the factor theorem f(-4), f(2) and f(5) I'd try f(-4) is all... if you get any there f(-4) = 0 then that is a possibility, then try f(2) if f(2) = 0 try f(5) the equation where f(-4) = 0, f(2) = 0 and f(5) = 0 is a solution

OpenStudy (mathstudent55):

If a, b, and c are roots of a polynomial, then the polynomial is (x - a)(x - b)(x - c) = 0

OpenStudy (mathstudent55):

Replace a, b, and c with -4, 2, and 5 and multiply out the three binomials.

OpenStudy (anonymous):

the answer is c?

OpenStudy (mathstudent55):

It's d.

OpenStudy (anonymous):

what!!!!!!!!!!

OpenStudy (mathstudent55):

(x - (-4))(x - 2)(x - 5) = (x + 4)(x - 2)(x - 5) = (x^2 + 2x - 8)(x - 5) = x^3 - 3x^2 - 18x + 40

OpenStudy (mathstudent55):

Notice the binomials are (x - a)(x - b)(x - c) (x - (-4)) = x + 4 (x - 2) = x - 2 (x - 5) = x - 5